# The dimension and Bose distance of certain primitive BCH codes

### v2

@@ IV The dimension of $\mathcal{C}_{(q,m,\delta)}$ for $2\leq\delta\leq q^{\lfloor({2m-1})/{3}\rfloor+1}$
$\sum\limits_{k=1}^{k_{\delta}-1}({k}-1)q^{2{k}-3}(q-1)^{3}(q+1)=\frac{\left[q-(k_{\delta}-1)q^{2k_{\delta}-3}+(k_{\delta}-2)q^{2k_{\delta}-1}\right](q-1)}{q+1}.$
@@
$\left|\left[q^{h+k_{\delta}},\delta-1\right]\cap\mathcal{S}\right|$ $=$ $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{S}\right|+\left|\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta q^{\ell},\delta-1\right]\cap\mathcal{S}\right|$ (47)
$=$ $\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)+N(\mu(\delta)+1).$
@@
$\left|\left[1,\delta-1\right]\cap\mathcal{S}\right|$ $=$ $\left|\left[1,q^{h+1}-1\right]\cap\mathcal{S}\right|+\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{S}\right|+\left|\left[q^{h+k_{\delta}},\delta-1\right]\cap\mathcal{S}\right|$ (48)
$=$ $(k_{\delta}-1)q^{2k_{\delta}-3}(q-1)^{2}+N(\mu(\delta)+1)+\sum\limits_{i=k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1).$
@@
$\left|\left[1,\delta-1\right]\cap(\mathcal{S}\cup\mathcal{H})\right|$ $=$ $\frac{1}{2}(\delta_{h}-1)\delta_{h}+N(\widetilde{\mu}(\delta)+1).$ (55)
@@
$\left|\left[q^{h+k},q^{h+k+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|=\sum\limits_{i=-k}^{k}|\mathcal{B}_{k}(i)|=(k-\frac{1}{2})q^{2k-2}(q-1)^{3}(q+1)+(q-1)^{2}(\frac{1}{2}q^{k-1}+q^{2k}).$
@@
$\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$ $=$ $\sum\limits_{k=1}^{k_{\delta}-1}\left|\left[q^{h+k},q^{h+k+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$
@@
$\left|\left[q^{h+k_{\delta}},\delta-1\right]\cap(\mathcal{S}\cup\mathcal{H})\right|=\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}N(tq^{2i}+1)+N(\widetilde{\mu}(\delta)+1).$ (60)
@@
$\left|\left[1,\delta-1\right]\cap\left(\mathcal{S}\cup\mathcal{H}\right)\right|$ $=$ $\left|\left[1,q^{h+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|+\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|+\left|\left[q^{h+k_{\delta}},\delta-1\right]\cap(\mathcal{S}\cup\mathcal{H})\right|$
$=$ $(k_{\delta}-\frac{1}{2})q^{2k_{\delta}-2}(q-1)^{2}+\frac{1}{2}q^{k_{\delta}-1}(q-1)+N(\widetilde{\mu}(\delta)+1)+\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}N(tq^{2i}+1)$
@@
$\begin{split}|[q^{h+1},q^{h+k_{\delta}})\cap\mathcal{H}|=\sum\limits_{k=1}^{k_{\delta}-1}|[q^{h+k},q^{h+{k}+1})\cap\mathcal{H}|=(q^{k_{\delta}-1}-1)(q-1).\end{split}$
@@
$\left|\left[1,\delta-1\right]\cap\mathcal{H}\right|$ $=$ $|[1,q^{h+k_{\delta}})\cap\mathcal{H}|+\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\right|+\left|\left[\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{H}\right|$
$=$ $N(\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h})+\tau(\delta).$
@@
Note that $\delta\leq q^{\lfloor({2m-1})/{3}\rfloor+1}\leq q^{m-\lfloor{m}/{4}\rfloor}$ for even $m$. Thus, by Theorem 1, we have $C_{a}=m$ or $|C_{a}|={m}/{2}$ for any integer $a\in[1,\delta-1]$. Consequently, we conclude that the equality in (31) holds. This completes the proof of Theorem 3. ∎
@@
$\mathrm{dim}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}n-mN(\delta),&\hbox{if }\delta\leq q^{h},\\ n-m\left(N(\delta)-\frac{1}{2}\delta_{h}^{2}\right),&\hbox{if }\delta>q^{h}\hbox{ and }\delta_{h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ n-m\left[N(\delta)-\frac{1}{2}(\delta_{h}-1)^{2}-\delta_{0}\right],&\hbox{if }\delta>q^{h}\hbox{ and }\delta_{h}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
@@
$\mathrm{dim}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}n-mN(\delta),&\hbox{if }\delta\leq q^{h+1},\\ n-m\left[N(\delta)-(q-1)\delta_{h+1}^{2}-(\delta_{h}-1)\delta_{h+1}-\delta_{0}\right],&\hbox{if }\delta>q^{h+1},\delta_{h}>0\hbox{ and }\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ n-m\left[N(\delta)-(q-1)\delta_{h+1}^{2}-\delta_{h}\delta_{h+1}\right],&\hbox{if }\delta>q^{h+1},\delta_{h}>0\hbox{ and }\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ n-m\left[N(\delta)-(q-1)(\delta_{h+1}^{2}-\delta_{h+1}+\delta_{1})-\delta_{0}\right],&\hbox{if }\delta>q^{h+1},\delta_{h}=0\hbox{ and }\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ n-m\left[N(\delta)-(q-1)\delta_{h+1}^{2}\right],&\hbox{if }\delta>q^{h+1},\delta_{h}=0\hbox{ and }\delta_{h+1}q\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
@@
Case 2. If $\delta>q^{h+1}$, we have $k_{\delta}=0$, then we distinguish the following subcases:
@@
Noticing that $\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{0}$ if $\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$, we can substitute these terms into equation (27) to obtain
$f(\delta)=\begin{cases}(q-1)\delta_{h+1}^{2}+(\delta_{h}-1)\delta_{h+1}+\delta_{0},&\hbox{if }\delta_{h}>0\hbox{ and }\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ (q-1)\delta_{h+1}^{2}+\delta_{h}\delta_{h+1},&\hbox{if }\delta_{h}>0\hbox{ and}\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
@@
Noting that $\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{1}q+\delta_{0}$ if $\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$, we can substitute these terms into equation (27) to derive
$f(\delta)=\begin{cases}(q-1)(\delta_{h+1}^{2}-\delta_{h+1}+\delta_{1})+\delta_{0},&\hbox{if }\delta_{h}=0\hbox{ and }\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ (q-1)\delta_{h+1}^{2},&\hbox{if }\delta_{h}=0\hbox{ and }\delta_{h+1}q\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
@@
V The Dimension of $\mathcal{C}_{(q,m,\delta,b)}$ for $\delta+b\leq q^{\lfloor({2m-1})/{3}\rfloor+1}+1$
@@
Let $m\geq 4$ be an integer, and let $\delta$ and $b$ be positive integers such that $\delta+b\leq q^{\lfloor({2m-1})/{3}\rfloor+1}+1.$
@@
    $\mathrm{dim}(\mathcal{C}_{(q,m,\delta,b)})$ $=$ $n+m\left[N(b)-f(b)\right]-m\left[N(b+\delta-1)-f(b+\delta-1)\right].$ (64)
@@
    $\mathrm{dim}(\mathcal{C}_{(q,m,\delta,b)})=n+m\left[N(b)-\widetilde{f}(b)\right]-m\left[N(b+\delta-1)-\widetilde{f}(b+\delta-1)\right]-\frac{m}{2}[g(b+\delta-1)-g(b)].$ (65)
@@
If $m$ is odd, recalling equation (49) in the proof of Theorem 3 and noting that $\delta+b\leq q^{\lfloor({2m-1})/{3}\rfloor+1}+1,$ we can conclude that
$\left|\mathcal{L}(b,\delta)\right|=\left|\mathcal{L}(1,b+\delta-1)\right|-\left|\mathcal{L}(1,b)\right|=\left[N(b+\delta-1)-f(b+\delta-1)\right]-\left[N(b)-f(b)\right].$
In addition, we can apply Theorem 1 to conclude that $|C_{a}|=m$ for all $a\in[b,b+\delta-2]$. Consequently, the equality in (64) follows.
If $m$ is even, since $\delta+b\leq q^{\lfloor({2m-1})/{3}\rfloor+1}+1$, we can apply Theorem 1 to conclude that $|C_{a}|=m$ or ${m}/{2}$ for any integer $a\in[b,b+\delta+2]$. Recalling equation (63), we can obtain
$\left|\mathcal{L}_{m}(b,\delta)\right|=\left|\mathcal{L}_{m}(1,b+\delta-1)\right|-\left|\mathcal{L}_{m}(1,b)\right|=\left[N(b+\delta-1)-\widetilde{f}(b+\delta-1)\right]-\left[N(b)-\widetilde{f}(b)\right]$
and
$\left|\mathcal{L}_{\frac{m}{2}}(b,\delta)\right|=\left|\mathcal{L}_{\frac{m}{2}}(1,b+\delta-1)\right|-\left|\mathcal{L}_{\frac{m}{2}}(1,b)\right|=g(b+\delta-1)-g(b).$
Then we can obtain equation (65). ∎
Remark 7.
One can easily verify that when $b=1$,
$N(b)=0,\quad f(b)=0,\quad\widetilde{f}(b)=0,\quad\text{and}\quad g(b)=0.$
Therefore, the equalities in (64) and (65) reduce to the equalities in (31) and (30), respectively. In other words, Theorem 4 includes the result in Theorem 3 as a special case.
VI The Bose distance of $\mathcal{C}_{(q,m,\delta,b)}$ for $\delta+b\leq q^{\lfloor({2m-1})/{3}\rfloor+1}+1$
Recall the relationship between the Bose distance of narrow-sense BCH codes and those that are not necessarily narrow-sense, as presented in equation (4). To obtain the Bose distance of $C_{(q,m,\delta,b)}$ for positive integers $b$ and $\delta$ with $\delta+b\leq q^{\lfloor({2m-1})/{3}\rfloor+1}+1$, it is sufficient to determine the Bose distance of $C_{(q,m,\delta)}$ for $\delta\in[2,q^{\lfloor({2m-1})/{3}\rfloor+1}].$ When $\delta=q^{\lfloor(2m-1)/3\rfloor+1}$, the Bose distance was given by [15, Theorem 13]. In this section, we determine the dimension of $C_{(q,m,\delta)}$ for $\delta\in[2,q^{\lfloor({2m-1})/{3}\rfloor+1}).$
@@
$d_{B}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}\delta,&\hbox{ if }q\nmid\delta,\\ \delta+1,&\hbox{ if }q\mid\delta.\\ \end{cases}$
@@
Note that $m-h=h$ when $m$ is odd, and $m-h=h+1$ when $m$ is even. We can apply Lemma 2 to conclude that $\delta$ is the smallest coset leader in $[\delta,n]$ if $q\nmid\delta$, and $\delta+1$ is the smallest coset leader in $[\delta,n]$ if $q\mid\delta.$ Then the result follows. ∎
@@
Suppose that $m\geq 5$ is an odd integer and $\delta\in\left[q^{h+1},q^{\lfloor({2m-1})/{3}\rfloor+1}\right)$ is an integer. Let $j_{\delta}$ be the integer such that $q^{h+j_{\delta}}\leq\delta<q^{h+j_{\delta}+1}$. Let $\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$ denote the $q$-adic expansion of $\delta$, and let $r_{\delta}$ be the smallest integer in $[-j_{\delta}+1,j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$. Define $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}.$
@@
    $d_{B}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}\delta,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)},\\ \hat{\delta}+1,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\\ &\hbox{ and }\delta_{h-r_{\delta}+1}\neq q-1,\\ \hat{\delta}+2,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\\ &\hbox{ and }\delta_{h-r_{\delta}+1}=q-1.\\ \end{cases}$ (66)
@@
    $d_{B}(\mathcal{C}_{(q,m,\delta)})=\begin{cases}\delta+1,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\geq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)},\\ \hat{\delta}+1,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}<\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\\ &\hbox{ and }\delta_{h-r_{\delta}+1}\neq q-1,\\ \hat{\delta}+2,&\hbox{ if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}<\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\\ &\hbox{ and }\delta_{h-r_{\delta}+1}=q-1.\end{cases}$ (67)
@@
$V(\hat{\delta})=(\mathbf{0}_{h-j_{\delta}},\delta_{h+j_{\delta}},\ldots,\delta_{h+r_{\delta}},\mathbf{0}_{h-j_{\delta}},\delta_{h+j_{\delta}},\ldots,\delta_{h-r_{\delta}+1}).$

### v3

@@ II Preliminaries
$\sum\limits_{k=1}^{k_{\delta}-1}({k}-1)q^{2{k}-3}(q-1)^{3}(q+1)\\ =\frac{\left[q-(k_{\delta}-1)q^{2k_{\delta}-3}+(k_{\delta}-2)q^{2k_{\delta}-1}\right](q-1)}{q+1}.$
@@
$\left|\left[q^{h+k_{\delta}},\delta-1\right]\cap\mathcal{S}\right|$ $=$ $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{S}\right|$ (45)
$+\>\left|\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta q^{\ell},\delta-1\right]\cap\mathcal{S}\right|$
$=$ $\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)$
$+\>N(\mu(\delta)+1).$
@@
$\left|\left[1,\delta-1\right]\cap\mathcal{S}\right|$ $=$ $\left|\left[1,q^{h+1}-1\right]\cap\mathcal{S}\right|+\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{S}\right|$ (46)
$+\>\left|\left[q^{h+k_{\delta}},\delta-1\right]\cap\mathcal{S}\right|$
$=$ $(k_{\delta}-1)q^{2k_{\delta}-3}(q-1)^{2}+N(\mu(\delta)+1)$
$+\>\sum\limits_{i=k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1).$
@@
$\left|\left[1,\delta-1\right]\cap(\mathcal{S}\cup\mathcal{H})\right|$ $=$ $\frac{1}{2}(\delta_{h}-1)\delta_{h}+N(\widetilde{\mu}(\delta)+1).\quad$ (53)
@@
${\left|\left[q^{h+k},q^{h+k+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|}$ $=$ $\sum\limits_{i=-k}^{k}|\mathcal{B}_{k}(i)|$
$=$ $(k-\frac{1}{2})q^{2k-2}(q-1)^{3}(q+1)$
$+\>(q-1)^{2}(\frac{1}{2}q^{k-1}+q^{2k}).$
@@
[…]
$=$ $\sum\limits_{k=1}^{k_{\delta}-1}\left|\left[q^{h+k},q^{h+k+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$
@@
$\left|\left[q^{h+k_{\delta}},\delta-1\right]\cap(\mathcal{S}\cup\mathcal{H})\right|$ $=$ $\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}N(tq^{2i}+1)$ (58)
$+\>N(\widetilde{\mu}(\delta)+1).$
@@
$\left|\left[1,\delta-1\right]\cap\left(\mathcal{S}\cup\mathcal{H}\right)\right|$ $=$ $\left|\left[1,q^{h+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$
$+\>\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$
$+\>\left|\left[q^{h+k_{\delta}},\delta-1\right]\cap(\mathcal{S}\cup\mathcal{H})\right|$
$=$ $(k_{\delta}-\frac{1}{2})q^{2k_{\delta}-2}(q-1)^{2}$
$+\>\frac{1}{2}q^{k_{\delta}-1}(q-1)+N(\widetilde{\mu}(\delta)+1)$
$+\>\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}N(tq^{2i}+1)$
@@
$\begin{split}\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{H}\right|&=\sum\limits_{k=1}^{k_{\delta}-1}|[q^{h+k},q^{h+{k}+1})\cap\mathcal{H}|\\ &=(q^{k_{\delta}-1}-1)(q-1).\end{split}$
@@
$\left|\left[1,\delta\!-\!1\right]\cap\mathcal{H}\right|$ $=$ $|[1,q^{h+k_{\delta}})\cap\mathcal{H}|+\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\right|$
$+\>\left|\left[\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{H}\right|$
$=$ $N\left(\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\right)+\tau(\delta).$
@@
Note that $\delta\leq q^{\lfloor({2m-1})/{3}\rfloor+1}\leq q^{m-\lfloor{m}/{3}\rfloor}$ for even $m$. Thus, by Theorem 1, we have $C_{a}=m$ or $|C_{a}|={m}/{2}$ for any integer $a\in[1,\delta-1]$. Consequently, we conclude that the equality in (29) holds. This completes the proof of Theorem 3. ∎
@@
[…]
@@
[…]
@@
Case 2. If $\delta>q^{h+1}$, we have $k_{\delta}=1$, then we distinguish the following subcases:
@@
Noticing that $\mu(\delta)=\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{0}$ if $\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$, we can substitute these terms into equation (25) to obtain
$f(\delta)=\begin{cases}(q-1)\delta_{h+1}^{2}+(\delta_{h}-1)\delta_{h+1}+\delta_{0},\\ \quad\quad\hbox{if }\delta_{h}>0\hbox{ and }\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ (q-1)\delta_{h+1}^{2}+\delta_{h}\delta_{h+1},\\ \quad\quad\hbox{if }\delta_{h}>0\hbox{ and }\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
@@
Noting that $\mu(\delta)=\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{1}q+\delta_{0}$ if $\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$, we can substitute these terms into equation (25) to derive
$f(\delta)=\begin{cases}(q-1)\delta_{h+1}^{2},\hbox{ if }\delta_{h}=0\hbox{ and }\delta_{h+1}q\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ (q-1)(\delta_{h+1}^{2}-\delta_{h+1}+\delta_{1})+\delta_{0},\\ \quad\quad\hbox{if }\delta_{h}=0\hbox{ and }\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
@@
[…]
@@
Let $m\geq 4$, and let $b$ and $\delta$ be two positive integers such that $b\geq 2$ and $b(q-1)+1\leq\delta\leq q^{\lfloor({2m-1})/{3}\rfloor+1}+1-b$.
@@
    $\mathrm{dim}(\mathcal{C}_{(q,m,\delta,b)})=n-m\left[N(b+\delta-1)-f(b+\delta-1)\right].$
@@
    $\begin{split}\mathrm{dim}(\mathcal{C}_{(q,m,\delta,b)})=&n-m\left[N(b+\delta-1)-\widetilde{f}(b+\delta\!-\!1)\right]\\ &-\frac{m}{2}g(b+\delta\!-\!1).\end{split}$
@@
Since $\delta\geq b(q-1)+1,$ for each integer $x\in[1,b-1]$, there exists $i$ such that $b\leq xq^{i}\leq bq-1\leq b+\delta-2$. Therefore,
$C_{x}\subseteq\bigcup\limits_{a=b}^{b+\delta-2}C_{a}\quad\hbox{for all }x\in[1,b-1].$
It follows that
$\bigcup\limits_{a=1}^{b+\delta-2}C_{a}=\bigcup\limits_{a=b}^{b+\delta-2}C_{a},$
which is equivalent to $\mathcal{C}_{(q,m,\delta,b)}=\mathcal{C}_{(q,m,b+\delta-1)}.$ Noticing that $b+\delta-1\leq q^{\lfloor({2m-1})/{3}\rfloor+1}$, we can apply Theorem 3 to obtain the result. ∎
[…]
When $\delta=q^{\lfloor(2m-1)/3\rfloor+1}$, the Bose distance was given by [15, Theorem 13]. In this section, we determine the dimension of $C_{(q,m,\delta)}$ for $\delta\in[2,q^{\lfloor({2m-1})/{3}\rfloor+1}).$
@@
[…]
@@
Note that mh=hm-h=h when $m$ is odd, and mh=h+1m-h=h+1 when $m$ is even. We can apply Lemma 2 to conclude that $\delta$ is the smallest coset leader in $\delta,n]$ if $q\nmid\delta$, and $\delta+1$ is the smallest coset leader in $\delta,n]$ if $q\mid\delta.$ Then the result follows. ∎
@@
Suppose that $m\geq 5$ is an odd integer and $\delta\in\left[q^{h+1},q^{\lfloor({2m-1})/{3}\rfloor+1}\right)$ is an integer. Let $j_{\delta}$ be the integer such that $q^{h+j_{\delta}}\leq\delta<q^{h+j_{\delta}+1}$. Let $\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$ denote the $q$-adic expansion of $\delta$, and let $r_{\delta}$ be the smallest integer in $j_{\delta}+1,j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$. Define $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}.$
@@
[…]
@@
    $d_{B}(\mathcal{C}_{(q,m,\delta)})=\\ \begin{cases}\delta+1,&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\geq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)},\\ \hat{\delta}+1,&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}<\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\\ &\hbox{and }\delta_{h-r_{\delta}+1}\neq q-1,\\ \hat{\delta}+2,&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}<\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\\ &\hbox{and }\delta_{h-r_{\delta}+1}=q-1.\end{cases}$ (63)
@@
$V(\hat{\delta})=(\mathbf{0}_{h-j_{\delta}},\delta_{h+j_{\delta}},\ldots,\delta_{h+r_{\delta}},\\ \mathbf{0}_{h-j_{\delta}},\delta_{h+j_{\delta}},\ldots,\delta_{h-r_{\delta}+1}).$ (64)